OpenStudy (anonymous):

The ratio of the perimeters of two rectangles is 5 to 8. The perimeter of the larger rectangle is 72 inches. What is the perimeter of the smaller rectangle? A. 115.2 inches B. 45 inches C. 40 inches D. 16.4 inches

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@Conqueror

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

LOL CROSS MULTIPLY

OpenStudy (anonymous):

how would i set that up?

OpenStudy (anonymous):

\[\frac{8}{5}=\frac{x}{72}\] then cross multiply lol

OpenStudy (anonymous):

ok one sec

OpenStudy (igreen):

WHAT. YOU HYPOCRITE.

OpenStudy (anonymous):

just kidding never cross multiply

OpenStudy (anonymous):

\[\frac{8}{5}=\frac{x}{72}\iff x=\frac{72\times 8}{5}\]

OpenStudy (igreen):

\(\color{blue}{\text{Originally Posted by}}\) @satellite73 \[\frac{8}{5}=\frac{x}{72}\] then cross multiply lol \(\color{blue}{\text{End of Quote}}\) This is how you set it up. Cross multiplying gives us: 8 * 72 = 5 * x

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

avoid cross multiplying whenever possible why multiply and then undo it by division?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

8*72=576

OpenStudy (anonymous):

i wrote the answer above, all you need to do is compute it

OpenStudy (anonymous):

115.2